OpenStudy (anonymous):

integrate by parts sin^(-1)xdx

OpenStudy (turingtest):

try\[u=\sin^{-1}x\\dv=dx\]

OpenStudy (anonymous):

I tried that and I always get stuck once I plug everything into the equation uv-integral vdu

OpenStudy (turingtest):

what do you get when you plug in?

OpenStudy (anonymous):

\[\sin ^{-1}x^2-\int\limits((x)/(1-x^2))dx\]

OpenStudy (turingtest):

the integral is right, but how do you get sin^{-1}(x^2)

OpenStudy (anonymous):

sin^(-1)x*x that was my u*v.

OpenStudy (anonymous):

is it just xsin^(-1)x

OpenStudy (turingtest):

right, not for that integral, do a *simple*, very common u-sub u=(what's under the radical)

OpenStudy (turingtest):

now for*

OpenStudy (turingtest):

oh you forgot the radical

OpenStudy (turingtest):

\[u=\sin^{-1}x\implies du=?\]

OpenStudy (anonymous):

\[1/(\sqrt{1-x^2})\]

OpenStudy (turingtest):

so \[\large\int vdu=?\]

OpenStudy (anonymous):

\[(1/\sqrt{(1-x^2}))xdx\]

OpenStudy (anonymous):

this is where i get stuck

OpenStudy (turingtest):

good, so now it's a simple u-sub let u=(what's under the radical) then du=?

OpenStudy (anonymous):

\[(1/2)(1-x^2)^{-1/2}*(-2x)dx\] ??

OpenStudy (turingtest):

yes, now you can integrate it

OpenStudy (turingtest):

still stuck?

OpenStudy (anonymous):

it doesn't make sense to have du be that long of a derivative

OpenStudy (turingtest):

why not? simply substitute in what we used as our u-substitution in the second integral u=1-x^2 du=-2xdx

OpenStudy (turingtest):

what does that give when you substitute it into the integral?

OpenStudy (anonymous):

\[xsin ^{-1}x-\int\limits(du/\sqrt{u)}\] ?

OpenStudy (turingtest):

you dropped the -1/2

OpenStudy (anonymous):

oh wait its \[xsin ^{-1}x-\int\limits(-1/2)du/(\sqrt{u)}dx\]

OpenStudy (turingtest):

yes, except now you pulled a dx out from somewhere.... not sure why :P

OpenStudy (anonymous):

on my paper i don't have the dx, sorry.

OpenStudy (turingtest):

no worries, so we can pull out that constant, giving\[x\sin^{-1}x+\frac12\int {du\over \sqrt u}\]

OpenStudy (turingtest):

which you should be able to integrate

OpenStudy (anonymous):

I have that

OpenStudy (turingtest):

ok, so what's the problem?

OpenStudy (anonymous):

I know what du/u is, ln(u)+c, but not du/sqrt. u.

OpenStudy (turingtest):

try writing 1/sqrtx as a fractional exponent